OpenStudy (jazzyfa30):

Use the remainder theorem to find the value for each when: f(x) = x4 + 3x2 + 5 g(x) = x3 -3x2 +x -7 g(3)

OpenStudy (jazzyfa30):

@Loser66

OpenStudy (rsadhvika):

put x = 3 in g(x)

OpenStudy (rsadhvika):

g(x) = x3 -3x2 +x -7 g(3) = 3^3 - 3(3)^2 + 3 -7

OpenStudy (rsadhvika):

simplify

OpenStudy (loser66):

yes, rsadhvika is right. to find the remainder, you just plug yhe number into the function and calculate it . that's it

OpenStudy (loser66):

g(3) = 27 - 27 +3-7 =-4

OpenStudy (jazzyfa30):

i got that -4

OpenStudy (loser66):

ok that's ok

OpenStudy (loser66):

you don't understand why it is negative, right?

OpenStudy (jazzyfa30):

yes

OpenStudy (loser66):

ok take the simple example tell me what is the remainder when i calculate 9/4

OpenStudy (jazzyfa30):

2

OpenStudy (loser66):

remainder?

OpenStudy (jazzyfa30):

yes 2

OpenStudy (loser66):

9 divided by 4 the quotient is 2 and remainder is 1 because 2*4 +1 =9

OpenStudy (jazzyfa30):

oh ya my b

OpenStudy (loser66):